OpenStudy (anonymous):

What number is 32% of 75? Which of the following equations could be used to solve the problem? n = 0.32(75) 0.32n = 75 n = 75 ÷ 0.32 This is just a reassurance question It's n = 75 ÷ 0.32 Right?

OpenStudy (jack12):

.32*75

terenzreignz (terenzreignz):

32% of 75 translates into \[\Large \frac{32}{100}\times 75\] Which is \[\Large 0.32 \times 75\]

OpenStudy (anonymous):

ohh okay so n = 0.32(75)

terenzreignz (terenzreignz):

Yup :)